OpenStudy (anonymous):

(sqrtx - 3)/(sqrtx + 3) derivative pls

OpenStudy (anonymous):

\[y=\frac{\sqrt{x}-3}{\sqrt{x}+3}\]

OpenStudy (anonymous):

\[y'=\frac{(\sqrt{x}+3)\times \frac{1}{2\sqrt{x}}-(\sqrt{x}-3)\times \frac{1}{2\sqrt{x}}}{(\sqrt{x}+3)^2}\]

OpenStudy (anonymous):

by the quotient rule, then algebra

OpenStudy (anonymous):

probably easiest to clean this up by multiplying top and bottom by \[2\sqrt{x}\]

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

\[y=\frac{\sqrt{x}-3}{\sqrt{x}+3}=\frac{\sqrt{x}+3-6}{\sqrt{x}+3}=1-\frac{6}{\sqrt{x}+3}\]\[y \prime=\frac{-6(1/2\sqrt{x})}{(\sqrt{x}+3)^{2}}\]